# Nyström Kernel Stein Discrepancy

### v2

@@ 5 Experiments
We verify the viability of our proposed method, abbreviated as N-KSD in this section, by comparing its runtime and its power to existing methods: the quadratic time KSD [Liu et al., 2016, Chwialkowski et al., 2016], the linear-time goodness-of-fit test finite set Stein discrepancy (FSSD; Jitkrittum et al. 2017), RFF-based KSD approximations [Huggins and Mackey, 2018], and the linear-time goodness-of-fit test using random feature Stein discrepancy (L1 IMQ, L2 SechExp; Huggins and Mackey 2018). [^6] All the code replicating our experiments is available at https://github.com/FlopsKa/nystroem-ksd. For FSSD, we consider randomized test locations (FSSD-rand) and optimized test locations (FSSD-opt); the optimality is meant w.r.t. a power proxy detailed in the cited work [Jitkrittum et al., 2017]. For all competitors, we use the settings and implementations provided by the respective authors. We use the well-known Gaussian kernel $k(\mathbf{x},\mathbf{y})=\exp\left(-\gamma\left\|{\mathbf{x}-\mathbf{y}}\right\|_{\mathbb{R}^{d}}^{2}\right)$ ($\gamma>0$) with the median heuristic [Garreau et al., 2018], and the IMQ kernel $k(\mathbf{x},\mathbf{y})=\left(c^{2}+\left\|{\mathbf{x}-\mathbf{y}}\right\|_{\mathbb{R}^{d}}^{2}\right)^{\beta}$ [Gorham and Mackey, 2017], with the choices of $\beta<0$ and $c>0$ detailed in the respective experiment description. To approximate the null distribution of N-KSD, we perform a bootstrap with (15), setting $D=500$. To allow an easy comparison, our experiments replicate goodness-of-fit testing experiments from Chwialkowski et al. [2016], Jitkrittum et al. [2017] and Huggins and Mackey [2018]. We ran all experiments on a PC with Ubuntu 20.04, 124GB RAM, and 32 cores with 2GHz each.
Runtime. We set $m=4\sqrt{n}$ for N-KSD to match the settings in our other experiments. As per recommendation, we fix the number of test locations $J=10$ for L1 IMQ, L2 SechExp, and both FSSD methods. The data is randomly generated and has $d=10$ dimensions. The runtime, see Figure 1(a) for the average over 10 repetitions (the error bars indicate the estimated $95\%$ quantile), behaves as predicted by the complexity analysis. The proposed approach runs orders of magnitudes faster than the quadratic time KSD estimator (7). From $n=1500$, all (near-)linear-time approaches are faster (excluding FSSD-opt, which has a relatively large fixed cost). Still, N-KSD achieves competitive runtime results even for $n=5000$.
@@
Restricted Boltzmann machine (RBM). Similar to Liu and Wang [2016], Jitkrittum et al. [2017], we consider the case where the target $\mathbb{P}$ is the non-normalized density of an RBM; the samples ^$\hat{\mathbb{Q}}_{n}$ are obtained from the same RBM perturbed by independent Gaussian noise with variance $\sigma^{2}$. For $\sigma^{2}=0$, $H_{0}:\mathbb{Q}=\mathbb{P}$ holds, and for $\sigma^{2}>0$, implying that the alternative $H_{1}:\mathbb{Q}\neq\mathbb{P}$ holds, the goal is to detect that the $n=1000$ samples come from a forged RBM. For the IMQ kernel (L1 IMQ, N-KSD IMQ, KSD IMQ), we set $c=1$ and $\beta=-1/2$. We show the results in Figure 1(d), using $100$ repetitions to obtain the power. KSD with the IMQ and with the Gaussian kernel performs best. Our proposed Nyström-based method ($m=4\sqrt{n}$) nearly matches its performance with the IMQ kernel while requiring only a fraction of the runtime. Besides Cauchy RFF and Gauss RFF, all other approaches achieve less power for $\sigma\in\{0.02,0.04\}$.
@@ Appendix A Proofs
This section is dedicated to the proofs of our results in the main text. Lemma 1 is proved in Section A.1. We prove our main result (Theorem 2) in Section A.2; Corollary 1 is shown in Section A.3.
@@ A.1 Proof of Lemma 1
In (a) we used that $\left\|\cdot\right\|_{\mathcal{H}_{h_{p},m}}$ is inner product induced, (b) follows from the linearity of the inner product, (c) is implied by the reproducing property, (d) is by the definition the Gram matrix, in (e) we made use of the explicit form of $\bm{\alpha}$, the symmetry of Gram matrices, the property $\mathbf{K}_{h_{p},m,n}^{\mkern-1.5mu\mathsf{T}}=\mathbf{K}_{h_{p},n,m}$, and that the Moore-Penrose inverse satisfies $\mathbf{A}^{-}\mathbf{A}\mathbf{A}^{-}=\mathbf{A}^{-}$ for any matrix $\mathbf{A}$.
@@ A.2 Proof of Theorem 2
Contrasting the existing related work [Rudi et al., 2015, Chatalic et al., 2022, Sterge and Sriperumbudur, 2022, Kalinke and Szabó, 2023], we do not impose a boundedness assumption on the feature map. This relaxation leads to new technical difficulties that we resolve in the following. Still, we start our analysis from a known decomposition Chatalic et al. [2022, Lemma 4.1].
@@
$\stackrel{{\scriptstyle(e)}}{{\leq}}\underbrace{\left\|{\mu_{h_{p}}-\hat{\mu}_{h_{p}}}\right\|_{\mathcal{H}_{h_{p}}}}_{=:t_{1}}+\underbrace{\left\|{\left(I-P_{\mathcal{H}_{h_{p},m}}\right)C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}}_{=:t_{2}}\underbrace{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\left(\hat{\mu}_{h_{p}}-\frac{1}{m}\sum_{i=1}^{m}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\right)}\right\|_{\mathcal{H}_{h_{p}}}}_{=:t_{3}}.$ (A.5)
(a) is implied by (A.1) and (11); (b) follows from the reverse triangle inequality; ±$\pm\hat{\mu}_{h_{p}}$ and the triangle inequality yield (c); in (d), we use the distributive property and introduce $\frac{1}{m}\sum_{i=1}^{m}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\in\mathcal{H}_{h_{p},m}$ whose projection onto the orthogonal complement of $\mathcal{H}_{h_{p},m}$ is zero; in (e) $I=C_{\mathbb{Q},h_{p},\lambda}^{1/2}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}$ was introduced and we used the definition of the operator norm.
@@
  • Term $t_{1}$. The first term measures the deviation of an empirical mean $\hat{\mu}_{h_{p}}$ to its population counterpart $\mu_{h_{p}}$. To bound this deviation $\left\|{\hat{\mu}_{h_{p}}-\mu_{h_{p}}}\right\|_{\mathcal{H}_{h_{p}}}=\Big{\|}\frac{1}{n}\sum_{i=1}^{n}[h_{p}\left(\cdot,{\mathbf{x}_{i}}\right)-\underbrace{\mathbb{E}_{X\sim\mathbb{Q}}h_{p}\left(\cdot,{X}\right)}_{=0}]\Big{\|}_{\mathcal{H}_{h_{p}}}$, we will use the Bernstein inequality (Theorem D.4) with the $\eta_{i}:=h_{p}\left(\cdot,{\mathbf{x}_{i}}\right)\in\mathcal{H}_{h_{p}}$ $(i\in[n])$ centered random variables, by gaining control on the moments of $Y:=\left\|h_{p}(\cdot,X)\right\|_{\mathcal{H}_{h_{p}}}$. This is what we elaborate next.
    By Assumption 1 and Lemma B.3, $Y$ is sub-Gaussian; hence it is also sub-exponential (Lemma D.2(3)), and therefore (Lemma B.2) it satisfies the Bernstein condition
@@
    for any $p\geq 2$. Notice that $B=\left\|{Y}\right\|_{\psi_{1}}\stackrel{{\scriptstyle(a)}}{{\lesssim}}\left\|{Y}\right\|_{\psi_{2}}\stackrel{{\scriptstyle(b)}}{{\lesssim}}\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}$. (a) follows from Lemma D.2(3) and (b) is implied by Lemma B.3. As $\sigma\asymp B$, we also got that $\sigma\lesssim\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}$.
    Having obtained a bound on the moments, we can apply Bernstein’s inequality for separable Hilbert spaces (Yurinsky 1995; recalled in Theorem D.4) to the centered $\eta_{i}=h_{p}\left(\cdot,{\mathbf{x}_{i}}\right)\in\mathcal{H}_{h_{p}}$-s ($i\in[n]$), and obtain that for any $\delta\in(0,1)$ it holds that
    $\mathbb{Q}^{n}\Bigg{(}\underbrace{\left\|{\mu_{h_{p}}-\hat{\mu}_{h_{p}}}\right\|_{\mathcal{H}_{h_{p}}}}_{(=\left\|{\frac{1}{n}\sum_{i=1}^{n}\eta_{i}}\right\|_{\mathcal{H}_{h_{p}}})}\lesssim\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}\log(6/\delta)}{n}+\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)\log(6/\delta)}{n}}\Bigg{)}\geq 1-\delta/3.\quad\quad$ (A.7)
@@
  • Term $t_{2}$. Assume that $0<\lambda\leq\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}$. Then, we can handle the second term with Lemma B.1 and obtain that for any $\delta\in(0,1)$ it holds that
    $\left(\mathbb{Q}^{n}\otimes\Lambda^{m}\right)\left(\left\|{\left(I-P_{\mathcal{H}_{h_{p},m}}\right)C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}\lesssim\sqrt{\lambda}\right)\geq 1-\delta/3$ (A.8)
    provided that $m\gtrsim\max\left\{\frac{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}{\lambda},\log(3/\delta)\right\}$.
@@
    $t_{3}$ $=\Bigg{\|}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\left(\hat{\mu}_{h_{p}}-\frac{1}{m}\sum_{i=1}^{m}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\right)\Bigg{\|}_{\mathcal{H}_{h_{p}}}$ $=\Bigg{\|}\frac{1}{m}\sum_{i=1}^{m}\underbrace{\left[C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\left(h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)-\hat{\mu}_{h_{p}}\right)\right]}_{:=Y_{i}}\Bigg{\|}_{\mathcal{H}_{h_{p}}}$ (A.9)
@@
    $\max_{i\in[m]}\left\|{Y_{i}}\right\|_{\mathcal{H}_{h_{p}}}$ $=\max_{i\in[m]}\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\left(h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)-\hat{\mu}_{h_{p}}\right)}\right\|_{\mathcal{H}_{h_{p}}}$ (A.10)
    $\leq\max_{i\in[m]}\left[\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}(\cdot,\tilde{\mathbf{x}}_{i})}\right\|_{\mathcal{H}_{h_{p}}}+\frac{1}{n}\sum_{j=1}^{n}\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}(\cdot,\mathbf{x}_{j})}\right\|_{\mathcal{H}_{h_{p}}}\right]$ (A.11)
    $\leq 2\max_{i\in[n]}\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{\mathbf{x}_{i}}\right)}\right\|_{\mathcal{H}_{h_{p}}}=:K=K(\mathbf{x}_{1},\ldots,\mathbf{x}_{n}),$ (A.12)
    where we used the triangle inequality and homogeneity of the norm. An application of Theorem D.5 yields that, conditioned on the sample $(\mathbf{x}_{i})_{i=1}^{n}$, it holds that
@@
    Randomness in $\mathbf{x}_{i}$-s: Let $Z_{i}:=\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{\mathbf{x}_{i}}\right)}\right\|_{\mathcal{H}_{h_{p}}}$ ($i\in[n]$) with $(\mathbf{x}_{i})_{i=1}^{n}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{Q}$. By Assumption 1 and Lemma B.3, the $Z_{i}$-s are sub-Gaussian random variables. Hence, by Lemma B.5, with probability at least $61-\delta/$ 6, it holds that
@@
    By Lemma B.3, $\left\|{Z_{1}}\right\|_{\psi_{2}}^{2}\lesssim\operatorname{tr}\left(C_{\mathbb{Q},h_{p},\lambda}^{-1}C_{\mathbb{Q},h_{p}}\right)$. We have shown that
    $\mathbb{Q}^{n}\!\left(\left(\mathbf{x}_{i}\right)_{i=1}^{n}\,:\,K\lesssim\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p},\lambda}^{-1}C_{\mathbb{Q},h_{p}}\right)\log(12n/\delta)}\right)\geq 1-\frac{\delta}{6}.$ (A.15)
@@
    $A$ =$=\left\{\left(\left(\mathbf{x}_{i}\right)_{i=1}^{n},\left(i_{j}\right)_{j=1}^{m}\right)\;:\;t_{3}\lesssim\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p},\lambda}^{-1}C_{\mathbb{Q},h_{p}}\right)\log(12n/\delta)\log(12/\delta)}}{\sqrt{m}}\right\},$ (A.16)
    $B$ =$=\left\{\left(\mathbf{x}_{i}\right)_{i=1}^{n}\,:\,K\lesssim\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p},\lambda}^{-1}C_{\mathbb{Q},h_{p}}\right)\log(12n/\delta)}\right\},$ (A.17)

### v4

@@ 5 EXPERIMENTS
We verify the viability of our proposed method, abbreviated as N-KSD in this section, by comparing its runtime and its power to existing methods: the quadratic time KSD [Liu et al., 2016, Chwialkowski et al., 2016], the linear-time goodness-of-fit test finite set Stein discrepancy (FSSD; Jitkrittum et al. 2017b), RFF-based KSD approximations [Huggins and Mackey, 2018], and the linear-time goodness-of-fit test using random feature Stein discrepancy (L1 IMQ, L2 SechExp; Huggins and Mackey 2018). [^8] The code replicating our experiments is available at https://github.com/FlopsKa/nystroem-ksd. For FSSD, we consider randomized test locations (FSSD-rand) and optimized test locations (FSSD-opt); the optimality is meant w.r.t. a power proxy detailed in the cited work. For all competitors, we use the settings and implementations provided by the respective authors. We use the well-known Gaussian kernel $k(\mathbf{x},\mathbf{y})=\exp\left(-\gamma\left\|{\mathbf{x}-\mathbf{y}}\right\|_{\mathbb{R}^{d}}^{2}\right)$ ($\gamma>0$) with the median heuristic [Garreau et al., 2018], and the IMQ kernel $k(\mathbf{x},\mathbf{y})=\left(c^{2}+\left\|{\mathbf{x}-\mathbf{y}}\right\|_{\mathbb{R}^{d}}^{2}\right)^{-\beta}$ [Gorham and Mackey, 2017], with the choices of $\beta,c>0$ detailed in the respective experiment description. To approximate the null distribution of N-KSD, we perform a bootstrap with (17), setting $D=500$. To allow an easy comparison, our experiments replicate goodness-of-fit testing experiments from Chwialkowski et al. [2016], Jitkrittum et al. [2017b] and Huggins and Mackey [2018]. For additional results, we refer to Appendix D. We ran all experiments on a PC with Ubuntu 20.04, 124GB RAM, and 32 cores with 2GHz each.
Runtime. We set $m=4\sqrt{n}$ for N-KSD to match the settings in our other experiments. As per recommendation, we fix the number of test locations $J=10$ for L1 IMQ, L2 SechExp, Cauchy RFF, Gauss RFF, and both FSSD methods. The data is randomly generated with $d=10$ dimensions. We note that the dimensionality enters the complexity only through the kernel evaluation; the dependence is linear in our case. The runtime, see Figure 1(a) for the average over 10 repetitions (the error bars indicate the estimated $95\%$ quantile), behaves as predicted by the complexity analysis. The proposed approach runs orders of magnitudes faster than the quadratic time KSD estimator (10). From $n=1500$, all (near-)linear-time approaches are faster (excluding FSSD-opt, which has a relatively large fixed cost). Still, N-KSD achieves competitive runtime results even for $n=5000$.
@@
Restricted Boltzmann machine (RBM). Similar to Liu et al. [2016], Jitkrittum et al. [2017b], we consider the case where the target $\mathbb{P}$ is the non-normalized density of an RBM with $50$ visible and $40$ hidden dimensions; the samples ^$\hat{\mathbb{Q}}_{n}$ are obtained from the same RBM perturbed by independent Gaussian noise with variance $\sigma^{2}$. For $\sigma^{2}=0$, $H_{0}:\mathbb{Q}=\mathbb{P}$ holds, and for $\sigma^{2}>0$, implying that the alternative $H_{1}:\mathbb{Q}\neq\mathbb{P}$ holds, the goal is to detect that the $n=1000$ samples come from a forged RBM. For the IMQ kernel (L1 IMQ, N-KSD IMQ, KSD IMQ), we set $c=1$ and $\beta=-1/2$. We show the results in Figure 1(d), using $100$ repetitions to obtain the power. KSD with the IMQ and with the Gaussian kernel performs best. Our proposed Nyström-based method ($m=4\sqrt{n}$) nearly matches its performance with the IMQ kernel while requiring only a fraction of the runtime. Besides Cauchy RFF and Gauss RFF, all other approaches achieve less power for $\sigma\in\{0.02,0.04\}$.
@@
6 LIMITATIONS
Assumption 1, which underpins our main result (Theorem 2), can be difficult to verify in some cases. We refer to Example 2 for a case where the analytical verification is possible. The weaker assumption $\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}<\infty$ of Theorem 3 is usually easier to verify analytically, as we show in Example 3. We note that, as with all kernel-based tests, the choice of the kernel, corresponding to the setting of $\gamma$ for the Gaussian kernel (resp. the setting of $\beta,c$ for the IMQ kernel), has an impact on the power of the test. While optimizing kernel parameters is not the focus of this work, there exist methods in the literature to (approximately) achieve this goal [Jitkrittum et al., 2016, 2017a, 2017b, Liu et al., 2020, Schrab et al., 2022a, b, Hagrass et al., 2024a, b, 2025].
@@ Appendix A PROOFS
This section is dedicated to the proofs of our results in the main text. Lemma 1 is proved in Section A.1. We prove our main result (Theorem 2) in Section A.2; Corollary 1 is shown in Section A.3. The proof of Theorem 3 is in Section A.4.
@@ A.1 Proof of Lemma 1
In (a) we used that $\left\|\cdot\right\|_{\mathcal{H}_{h_{p},m}}$ is inner product induced, (b) follows from the linearity of the inner product, (c) is implied by the reproducing property, (d) is by the definition of the Gram matrix, in (e) we made use of the explicit form of $\bm{\alpha}$, the symmetry of Gram matrices, the property $\mathbf{K}_{h_{p},m,n}^{\mkern-1.5mu\mathsf{T}}=\mathbf{K}_{h_{p},n,m}$, and that the Moore-Penrose inverse satisfies $\mathbf{A}^{-}\mathbf{A}\mathbf{A}^{-}=\mathbf{A}^{-}$ for any matrix $\mathbf{A}$.
@@ A.2 Proof of Theorem 2
Contrasting the existing related work [Rudi et al., 2015, Chatalic et al., 2022, Sterge and Sriperumbudur, 2022, Kalinke and Szabó, 2023], we do not impose a boundedness assumption on the feature map. This relaxation leads to new technical difficulties that we resolve in the following. We start our analysis from a decomposition similar to Chatalic et al. [2022, Lemma 4.1]; the difference is that we introduce the centered covariance operator $C_{\mathbb{Q},\bar{h}_{p},\lambda}$ which allows us to handle both $\mathbb{P}=\mathbb{Q}$ and the challenging case of $\mathbb{P}\neq\mathbb{Q}$ in a unified fashion.
@@
$\stackrel{{\scriptstyle(e)}}{{\leq}}\underbrace{\left\|{\mu_{h_{p}}-\hat{\mu}_{h_{p}}}\right\|_{\mathcal{H}_{h_{p}}}}_{=:t_{1}}+\underbrace{\left\|{\left(I-P_{\mathcal{H}_{h_{p},m}}\right)C_{\mathbb{Q},\bar{h}_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}}_{=:t_{2}}\underbrace{\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(\hat{\mu}_{h_{p}}-\frac{1}{m}\sum_{i=1}^{m}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\right)}\right\|_{\mathcal{H}_{h_{p}}}}_{=:t_{3}}.$ (A.5)
(a) is implied by (A.1) and (13); (b) follows from the reverse triangle inequality; ±$\pm\hat{\mu}_{h_{p}}$ and the triangle inequality yield (c); in (d), we use the distributive property and introduce the term $\frac{1}{m}\sum_{i=1}^{m}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\in\mathcal{H}_{h_{p},m}$ whose projection onto the orthogonal complement of $\mathcal{H}_{h_{p},m}$ is zero; in (e) $I=C_{\mathbb{Q},\bar{h}_{p},\lambda}^{1/2}C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}$ was introduced and we used the definition of the operator norm.
@@
  • Term $t_{1}$. The first term measures the deviation of an empirical mean $\hat{\mu}_{h_{p}}$ to its population counterpart $\mu_{h_{p}}$. To bound this deviation $\left\|{\hat{\mu}_{h_{p}}-\mu_{h_{p}}}\right\|_{\mathcal{H}_{h_{p}}}=\left\|\frac{1}{n}\sum_{i=1}^{n}\bar{h}_{p}\left(\cdot,{\mathbf{x}_{i}}\right)\right\|_{\mathcal{H}_{h_{p}}}$, we will use the Bernstein inequality (Theorem C.4) with the $\eta_{i}:=\bar{h}_{p}\left(\cdot,{\mathbf{x}_{i}}\right)\in\mathcal{H}_{h_{p}}$ $(i\in[n])$ centered random variables, by gaining control on the moments of $Y:=\left\|\bar{h}_{p}\left(\cdot,{X}\right)\right\|_{\mathcal{H}_{h_{p}}}$. This is what we elaborate next.
    By Assumption 1 and Lemma B.3, $Y$ is sub-Gaussian; hence it is also sub-exponential (Lemma C.2(3)), and therefore (Lemma B.2) it satisfies the Bernstein condition
@@
    for any $p\geq 2$. Notice that $B=\left\|{Y}\right\|_{\psi_{1}}\stackrel{{\scriptstyle(a)}}{{\lesssim}}\left\|{Y}\right\|_{\psi_{2}}\stackrel{{\scriptstyle(b)}}{{\lesssim}}\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}$. (a) follows from Lemma C.2(3) and (b) is implied by Lemma B.3. As $\sigma\asymp B$, we also got that $\sigma\lesssim\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}$.
    Having obtained a bound on the moments, we can apply Bernstein’s inequality for separable Hilbert spaces (Yurinsky 1995; recalled in Theorem C.4) to the centered $\eta_{i}=\bar{h}_{p}\left(\cdot,{\mathbf{x}_{i}}\right)\in\mathcal{H}_{h_{p}}$-s ($i\in[n]$), and obtain that for any $\delta\in(0,1)$ it holds that
    $\mathbb{Q}^{n}\Bigg{(}\underbrace{\left\|{\mu_{h_{p}}-\hat{\mu}_{h_{p}}}\right\|_{\mathcal{H}_{h_{p}}}}_{(=\left\|{\frac{1}{n}\sum_{i=1}^{n}\eta_{i}}\right\|_{\mathcal{H}_{h_{p}}})}\lesssim\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}\log(6/\delta)}{n}+\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)\log(6/\delta)}{n}}\Bigg{)}\geq 1-\delta/3.\quad\quad$ (A.7)
@@
  • Term $t_{2}$. Assume that $0<\lambda\leq\left\|{C_{\mathbb{Q},\bar{h}_{p}}}\right\|_{\mathrm{op}}$. Then, we can handle the second term with Lemma B.1 and obtain that for any $\delta\in(0,1)$ it holds that
    $\left(\mathbb{Q}^{n}\otimes\Lambda^{m}\right)\left(\left\|{\left(I-P_{\mathcal{H}_{h_{p},m}}\right)C_{\mathbb{Q},\bar{h}_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}\lesssim\sqrt{\lambda}\right)\geq 1-\delta/3$ (A.8)
    provided that $m\gtrsim\max\left\{\frac{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}{\lambda},1\right\}\log\left(12/\delta\right)$.
@@
    $t_{3}$ $=\Bigg{\|}C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(\hat{\mu}_{h_{p}}-\frac{1}{m}\sum_{i=1}^{m}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\right)\Bigg{\|}_{\mathcal{H}_{h_{p}}}=\Bigg{\|}\frac{1}{m}\sum_{i=1}^{m}\underbrace{\left[C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)-\hat{\mu}_{h_{p}}\right)\right]}_{=:Y_{i}}\Bigg{\|}_{\mathcal{H}_{h_{p}}}$ (A.9)
@@
    $\max_{i\in[m]}\left\|{Y_{i}}\right\|_{\mathcal{H}_{h_{p}}}=\max_{i\in[m]}\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)-\hat{\mu}_{h_{p}}\right)}\right\|_{\mathcal{H}_{h_{p}}}$ (A.10)
    $=\max_{i\in[m]}\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)-\mathbb{E}_{X\sim\mathbb{Q}}h_{p}\left(\cdot,{X}\right)+\mathbb{E}_{X\sim\mathbb{Q}}h_{p}\left(\cdot,{X}\right)-\hat{\mu}_{h_{p}}\right)}\right\|_{\mathcal{H}_{h_{p}}}$ (A.11)
    $\leq\max_{i\in[m]}\Big{\|}C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\big{(}\underbrace{h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)-\mathbb{E}_{X\sim\mathbb{Q}}h_{p}\left(\cdot,{X}\right)}_{=\bar{h}_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)}\big{)}\Big{\|}_{\mathcal{H}_{h_{p}}}+\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(\hat{\mu}_{h_{p}}-\mathbb{E}_{X\sim\mathbb{Q}}h_{p}\left(\cdot,{X}\right)\right)}\right\|_{\mathcal{H}_{h_{p}}}$ (A.12)
    $\leq\max_{i\in[n]}\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\bar{h}_{p}\left(\cdot,{\mathbf{x}_{i}}\right)}\right\|_{\mathcal{H}_{h_{p}}}+\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(\hat{\mu}_{h_{p}}-\mathbb{E}_{X\sim\mathbb{Q}}h_{p}\left(\cdot,{X}\right)\right)}\right\|_{\mathcal{H}_{h_{p}}}$ (A.13)
    $\leq\max_{i\in[n]}\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\bar{h}_{p}\left(\cdot,{\mathbf{x}_{i}}\right)}\right\|_{\mathcal{H}_{h_{p}}}+\frac{1}{n}\sum_{i\in[n]}\Big{\|}{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\big{(}\underbrace{h_{p}\left(\cdot,{\mathbf{x}_{i}}\right)-\mathbb{E}_{X\sim\mathbb{Q}}h_{p}\left(\cdot,{X}\right)}_{=\bar{h}_{p}\left(\cdot,{\mathbf{x}_{i}}\right)}\big{)}}\Big{\|}_{\mathcal{H}_{h_{p}}}$ (A.14)
    $\leq 2\max_{i\in[n]}\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\bar{h}_{p}\left(\cdot,{\mathbf{x}_{i}}\right)}\right\|_{\mathcal{H}_{h_{p}}}=:K=K(\mathbf{x}_{1},\ldots,\mathbf{x}_{n}),$ (A.15)
    where we used that ±$\pm\mathbb{E}_{X\sim\mathbb{Q}}h_{p}\left(\cdot,{X}\right)=0$, the triangle inequality, and the homogeneity of the norm. An application of Theorem C.5 yields that, conditioned on the sample $(\mathbf{x}_{i})_{i=1}^{n}$, it holds that
@@
    Randomness in $\mathbf{x}_{i}$-s: Let $Z_{i}:=\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\bar{h}_{p}\left(\cdot,{\mathbf{x}_{i}}\right)}\right\|_{\mathcal{H}_{h_{p}}}$ ($i\in[n]$) with $(\mathbf{x}_{i})_{i=1}^{n}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{Q}$. By Assumption 1 and Lemma B.3, the $Z_{i}$-s are sub-Gaussian random variables. Hence, by Lemma B.5, with probability at least $61-\delta/$ 6, it holds that
@@
    By Lemma B.3, $\left\|{Z_{1}}\right\|_{\psi_{2}}^{2}\lesssim\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1}C_{\mathbb{Q},\bar{h}_{p}}\right)$. We have shown that
    $\mathbb{Q}^{n}\!\left(\left(\mathbf{x}_{i}\right)_{i=1}^{n}\,:\,K\lesssim\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1}C_{\mathbb{Q},\bar{h}_{p}}\right)\log(12n/\delta)}\right)\geq 1-\frac{\delta}{6}.$ (A.18)
@@
    $A$ =$=\left\{\left(\left(\mathbf{x}_{i}\right)_{i=1}^{n},\left(i_{j}\right)_{j=1}^{m}\right)\;:\;t_{3}\lesssim\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1}C_{\mathbb{Q},\bar{h}_{p}}\right)\log(12n/\delta)\log(12/\delta)}}{\sqrt{m}}\right\},$ (A.19)
    $B$ =$=\left\{\left(\mathbf{x}_{i}\right)_{i=1}^{n}\,:\,K\lesssim\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1}C_{\mathbb{Q},\bar{h}_{p}}\right)\log(12n/\delta)}\right\},$ (A.20)